graph the equation 9y = x - 4

Answer by algebrahouse.com(1659)

Slope-intercept form is y = mx + b
m is the slope {rise over run}
b is the y-intercept {where the line crosses the y-axis}

9y = x - 4
y = (1/9)x - 4/9 {divided both sides by 9, to get into slope-intercept form}
slope = 1/9
y-intercept = -4/9

To graph:
- put a point on -4/9 on the y-axis
- from there, go up 1 and to the right 9 and put another point
- draw a line through the two points

For more help from me, visit: www.algebrahouse.com